OpenStudy (anonymous):

Help Please! (: A box with a square top and bottom and volume 20 cubic feet is to be constructed from two different materials. Exotic hardwood coasting 5 dollars per feet squared will be used for the top and bottom. Find the dimensions of the box that minimizes the cost of the materials.

zepdrix (zepdrix):

Since the base is square, two of the dimensions we can label the same: \[\Large\sf V=x\cdot x\cdot y\]They told us the volume is 20 cubic feet,\[\Large\sf 20=x^2 y\] And then we have this cost per square feet.... and we need to minimize cost.. hmmmm I'm having trouble coming up with another equation.. grr

OpenStudy (anonymous):

Wouldn't we have to take the derivative of the volume formula of a rectangle/cube? @zepdrix

zepdrix (zepdrix):

The derivative of volume will give us a formula for surface area. But that's the surface area of the entire cube. We only want the surface area of the top and bottom right? I dunno.. this question is really confusing..

zepdrix (zepdrix):

It just doesn't make sense... If we're trying to minimize cost, we would just make the top and bottom as tiny as possible because they didn't give us a price for the sides.

OpenStudy (anonymous):

how would we set this up? right sum? left sum? @zepdrix

zepdrix (zepdrix):

I don't know :( I don't understand this problem..

OpenStudy (anonymous):

@Loser66 @mathstudent55 Help?